OpenStudy (jessiegonzales):

MEDAL!!!!!!!!! 4. Is the following always, sometimes, or never true? 14+3x-7=7x+7-4x A. always B. sometimes C. never

OpenStudy (jessiegonzales):

@zepdrix

zepdrix (zepdrix):

Attempt to solve for x. If you end up with `x=#`, then it is `sometimes true`. This is telling us that the equation is true for ONE SPECIFIC X VALUE. If you end up with `5=6`, or something like that, then we have a false statement, 5 is never equal to 6, `never true`. If you end up with `3=3`, where both sides match, then your equation holds true for every x value, `always true`.

zepdrix (zepdrix):

Before adding or subtracting from both sides or anything like that, let's try to combine like terms first. \[\Large\rm \color{royalblue}{14}+3x\color{royalblue}{-7}=\color{orangered}{7x}+7\color{orangered}{-4x}\]

OpenStudy (jessiegonzales):

-7=3x=3x+7

zepdrix (zepdrix):

you meant to write this? -7+3x=3x+7 Hmm no we shouldn't get a -7 on the left side. 14-7 = ?

OpenStudy (jessiegonzales):

the result is A always

OpenStudy (jessiegonzales):

7

zepdrix (zepdrix):

You got the same thing on both sides? 7+3x = 3x+7 mmm ok good, that's correct! :)